OpenStudy (anonymous):

Solve 5c - c + 10 = 34. -6 6 11 -11

OpenStudy (dobby1):

ok I am going to explain this step by step I am not going to just give you the answer

OpenStudy (dobby1):

@kbbbb are you still there

OpenStudy (anonymous):

Okay thanks

OpenStudy (dobby1):

Ok the fist step is to combine like terms for example if I had b + 2b - 7 = 100 I would combine b and 2b giving me 3b but if it was b - 2b - 7 = 100 I would combine b and 2b giving me -b. So what would happen when you combine 5c and c. 5c - c + 10 = 34.

OpenStudy (dobby1):

@kbbbb

OpenStudy (dobby1):

you still there

OpenStudy (anonymous):

Yeah

OpenStudy (dobby1):

what would happen when you combine 5c with -c

OpenStudy (dobby1):

think of it as 5c - c

OpenStudy (dobby1):

come on I am not just going to give you the answer I am going to teach you how to do this.

OpenStudy (dobby1):

@kbbb what is 5c-c

OpenStudy (anonymous):

5?

OpenStudy (dobby1):

no 4c because c stands for 1c 5-1=4

OpenStudy (dobby1):

4c + 10 = 34

OpenStudy (anonymous):

Ohhhhhhhh

OpenStudy (anonymous):

So is c just a missing number?

OpenStudy (dobby1):

no it is 1 times whatever c equals same for 5c it is just 5 times whatever c equals

OpenStudy (anonymous):

Oh okay

OpenStudy (dobby1):

the next step is to get rid of the 10

OpenStudy (dobby1):

you do that by subtracting it from both sides

OpenStudy (anonymous):

Okay

OpenStudy (dobby1):

what do you get?

OpenStudy (dobby1):

Hint: when you subtract 10 from both sides it does not affect the 4c in any way

OpenStudy (dobby1):

yes 4c=24 now you have one more step and that is to divide each side by 4 this does affect the 4c by eliminating the 4

OpenStudy (anonymous):

What is the full problem right now..? I'm kind of confused of what the problem is supposed to look like..

OpenStudy (dobby1):

so c=?

OpenStudy (dobby1):

the problem looks like 4c=24 divide each side by 4 to get rid of the 4 and get the c by its self

OpenStudy (anonymous):

So 6?

OpenStudy (dobby1):

yes 6 can i have a medal

OpenStudy (dobby1):

thx